Sheldon Company manufactures only one product and uses a standard cost system. During the past month, the manufacturing operations had the following variances: Direct labor rate variance = $29,000 Favorable. Direct labor efficiency variance = $48,000 Unfavorable. Sheldon allows 4.60 standard direct labor hours per unit produced, and its standard direct labor hourly rate is $46. During the month, the company used 23.00% more direct labor hours than the standard allowed. What was the standard labor cost of units produced for the month? $48,000 $29,000 $227,696 $256,696 $208,696 Ventura uses a just-in-time (JIT) manufacturing system for all of its materials, components, and products. The master budget of the company for June called for use of 10,400 square feet of materials, while the flexible budget for the actual output of the month had 10,000 square feet of materials at a standard cost of 59.10 per square foot. Company records show that the actual price paid for the materials used in June was $9.08 per square foot, and that the direct materials purchase-price variance for the month was $980. The materials flexible budegt variance for June was $354,420 unfavorable 353,920 unfavorable 354,900 unfavorable 5355,400 unfavorable $980 favorable
Answer #1

Solution 1:

Standard rate of labor = $46 per hour

Let standard hours allowed = X hours

Actual direct labor hours = X * 123% = 1.23 X hours

Direct labor efficiency variance = $48,000 U

(SH - AH) * SR = -$48,000

(X - 1.23X) * $46 = -$48,000

X = 4536.86 hours

Standard direct labor cost = Standard hours allowed * standard rate = 4536.86 * $46 = $208,696

Hence last option is correct.

Solution 2:

Material purchase price variance = $980 F

(SP - AP) * AQ = $980

($9.10 - $9.08) * AQ = $980

AQ = 49000 Square foot

SQ for actual output = 10000 square feet

Material quantity variance = (SQ - AQ) * SR

= (10000 - 49000) * $9.10 = $354,900 U

Material flexible budget variance = Material price variance + Material quantity variance

= $980 F + $354,900 U = $353,920 U

Hence 2nd option is correct.